# From Age-Structured Trophic Networks to Applied Control : Stabilization and Harvesting Strategies for Non-Transitive Competition and the Dynamics of Mosquitoes

### v1

@@ 3.3. Toward generalization
(3.150) $\begin{aligned} u=&u^{*}+c_{N}\phi_{N}(z_{N-2})+\phi_{1}(z_{N-1})+\frac{c_{N}}{\lambda_{1}}\phi_{1}(z_{N-1})\phi_{N}(z_{N-2})-\phi_{2}(\eta_{2})-c_{1}\phi_{2}(\eta_{2})-\phi_{3}(z_{1})-\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})\\ &+c_{2}\phi_{3}(z_{1})+\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})+c_{N-1}\phi_{N}(z_{N-2})+\phi_{N+1}(z_{N-1})+\frac{c_{N-1}}{\lambda_{N+1}}\phi_{N+1}(z_{N-1})\phi_{N}(z_{N-2})\\ &-\frac{1}{\phi_{1}(z_{N})}\left(c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}\right)\\ &+\frac{\left(\phi_{1}(z_{N-1})-\theta\phi_{N+1}(z_{N-1})\right)}{\phi_{1}(z_{N})}\left(c_{N}\phi_{N}(z_{N-2})+\phi_{1}(z_{N-1})+\frac{c_{N}}{\lambda_{1}}\phi_{1}(z_{N-1})\phi_{N}(z_{N-2})\right)\\ &+\frac{\left(\phi_{1}(z_{N-1})-\theta\phi_{N+1}(z_{N-1})\right)}{\phi_{1}(z_{N})}\left(c_{1}\phi_{2}(\eta_{2})+\phi_{3}(z_{1})+\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})\right)\\ &+\frac{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)}{\phi_{1}(z_{N})}\left(c_{2}\phi_{3}(z_{1})+\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})\right)\\ &+\cdots+\frac{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)}{\phi_{1}(z_{N})}\left(c_{N-1}\phi_{N}(z_{N-2})+\phi_{N+1}(z_{N-1})+\frac{c_{N-1}}{\lambda_{N+1}}\phi_{N+1}(z_{N-1})\phi_{N}(z_{N-2})\right).\end{aligned}$
Moreover, one can construct a Lyapunov function $V_{N+1}$ of the form
(3.151) $V_{N+1}(\eta_{2},z_{1},\dots,z_{N})=V_{N}(\eta_{2},z_{1},\dots,z_{N-1})+(\theta\lambda_{N+1}-\lambda_{1})\left(e^{z_{N-1}}-1-z_{N-1}\right)+\lambda_{1}\left(e^{z_{N}}-1-z_{N}\right),$
such that
(3.152) $\begin{aligned} \dot{V}_{N+1}(\eta_{2},z_{1},\dots,z_{N})={}&\ \dot{V}_{N}(\eta_{2},z_{1},\dots,z_{N-1})\Big{|}_{\psi_{i}\equiv 0\ }\\ &\ +c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)},\quad(\psi_{i}\equiv 0).\end{aligned}$
and
(3.153) $\begin{aligned} \dot{V}_{N+1}(\eta_{2},z_{1},\dots,z_{N})={}&\ \dot{V}_{N}(\eta_{2},z_{1},\dots,z_{N-1})\Big{|}_{\psi_{i}\equiv 0\ }\\ &\ +c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}+\sum_{i=1}^{N+1}A_{i}\,\left(\hat{\phi}_{i}-\phi_{i}(\eta_{i})\right),\quad(\psi_{i}\neq 0).\end{aligned}$
@@
    By applying the Lyapunov functions defined in (3.61) and (3.121), we deduce the equivalent recursive form $N\mapsto N+1$ species
    (3.154) $V_{N+1}(\eta_{2},z_{1},\dots,z_{N})=V_{N}(\eta_{2},z_{1},\dots,z_{N-1})+(\theta\lambda_{N+1}-\lambda_{1})\left(e^{z_{N-1}}-1-z_{N-1}\right)+\lambda_{1}\left(e^{z_{N}}-1-z_{N}\right),$
    with $V_{N}$ defined by (3.147). Taking the derivative as follows :
    (3.155) $\dot{V}_{N+1}(\eta_{2},z_{1},\dots,z_{N})=\dot{V}_{N}(\eta_{2},z_{1},\dots,z_{N-1})+\theta\phi_{N+1}(z_{N-1})\dot{z}_{N-1}-\phi_{1}(z_{N-1})\dot{z}_{N-1}+\phi_{1}(z_{N})\dot{z}_{N}.$
    Since the generalization for $N\geq 3$ coincides with (3.148) for $\psi_{i}\equiv 0$ and with (3.149) for $\psi_{i}\neq 0$, the recursive relations linking $\dot{V}_{N}$ to $\dot{V}_{N+1}$ are
    (3.156) $\begin{aligned} \dot{V}_{N+1}(\eta_{2},z_{1},\dots,z_{N})={}&\ \dot{V}_{N}(\eta_{2},z_{1},\dots,z_{N-1})\Big{|}_{\ \psi_{i}\equiv 0\ }\\ &\ +c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)},\quad(\psi_{i}\equiv 0).\end{aligned}$
    and
    (3.157) $\begin{aligned} \dot{V}_{N+1}(\eta_{2},z_{1},\dots,z_{N})={}&\ \dot{V}_{N}(\eta_{2},z_{1},\dots,z_{N-1})\Big{|}_{\ \psi_{i}\neq 0\ }\\ &\ +c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}+\sum_{i=1}^{N+1}A_{i}\,\left(\hat{\phi}_{i}-\phi_{i}(\eta_{i})\right),\quad(\psi_{i}\neq 0).\end{aligned}$
    Now let us find the general control $u$ in both cases, i.e. $\psi_{i}\equiv 0$ and $\psi_{i}\neq 0.$
    Using (3.146), we get for $N+1$ species
    (3.158) $\begin{cases}z_{N}=\eta_{1}-\eta_{N+1}-\eta_{N}-\cdots-\eta_{3}+\eta_{2},\\ z_{N-1}=\eta_{N+1}-\eta_{N}-\eta_{N-1}-\cdots-\eta_{3}+\eta_{2}.\end{cases}$
Step 1 ($\psi_{i}\equiv 0$) : Let us find the control $u$ such that the following equality is satisfied
(3.159) $\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)\dot{z}_{N-1}+\phi_{1}(z_{N})\dot{z}_{N}=c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}$
We get from (3.146) for $N+1$ species
(3.160) $\begin{cases}\dot{z}_{N}=u^{*}-u-\phi_{2}+\phi_{1}+\phi_{N+1}+\phi_{N}+\cdots+\phi_{4}-\phi_{3},\\ \dot{z}_{N-1}=-\phi_{1}+\phi_{N+1}+\phi_{N}+\cdots+\phi_{4}-\phi_{3}\end{cases}$
Thus,
(3.161) $\begin{aligned} \left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)\dot{z}_{N-1}+\phi_{1}(z_{N})\dot{z}_{N}=&\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)\left(-\phi_{1}+\phi_{N+1}+\phi_{N}+\cdots+\phi_{4}-\phi_{3}\right)\\ &+\phi_{1}(z_{N})\left(-\phi_{2}+\phi_{1}+\phi_{N+1}+\phi_{N}+\cdots+\phi_{4}-\phi_{3}\right)+\phi_{1}(z_{N})(u^{*}-u)\end{aligned}$
(3.162) $\begin{aligned} \left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)\dot{z}_{N-1}+\phi_{1}(z_{N})\dot{z}_{N}=&\underbrace{\left(\phi_{1}(z_{N})-\theta\phi_{N+1}(z_{N-1})+\phi_{1}(z_{N-1})\right)}_{P_{1}}\phi_{1}\underbrace{-\phi_{1}(z_{N})}_{P_{2}}\phi_{2}\\ &+\underbrace{\left(\phi_{1}(z_{N-1})-\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N})\right)}_{P_{3}}\phi_{3}\\ &+\underbrace{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})+\phi_{1}(z_{N})\right)}_{P_{4}}\phi_{4}\\ &+\cdots\cdots+\underbrace{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})+\phi_{1}(z_{N})\right)}_{P_{N+1}}\phi_{N+1}+\phi_{1}(z_{N})(u^{*}-u)\end{aligned}$
From (3.159), we get
(3.163) $\begin{aligned} \phi_{1}(z_{N})(u^{*}-u)=&c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}\\ &-P_{1}\phi_{1}(\eta_{1})-P_{2}\phi_{2}(\eta_{2})-\cdots-P_{N+1}\phi_{N+1}(\eta_{N+1})\end{aligned}$
Finally, the control $u$ takes the following form
(3.164) $\begin{aligned} u=&u^{*}-\frac{1}{\phi_{1}(z_{N})}\left(c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}\right)\\ &+\frac{P_{1}}{\phi_{1}(z_{N})}\phi_{1}(\eta_{1})+\frac{P_{2}}{\phi_{1}(z_{N})}\phi_{2}(\eta_{2})+\frac{P_{3}}{\phi_{1}(z_{N})}\phi_{3}(\eta_{3})+\cdots+\frac{P_{N+1}}{\phi_{1}(z_{N})}\phi_{N+1}(\eta_{N+1})\end{aligned}$
satisfying (3.156). Using (3.94)-(3.131), equation
(3.165) $\begin{cases}\phi_{1}(\eta_{1})=c_{N}\phi_{N}(z_{N-2})+\phi_{1}(z_{N-1})+\frac{c_{N}}{\lambda_{1}}\phi_{1}(z_{N-1})\phi_{N}(z_{N-2}),\\ \\ \phi_{3}(\eta_{3})=c_{1}\phi_{2}(\eta_{2})+\phi_{3}(z_{1})+\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2}),\\ \\ \phi_{4}(\eta_{4})=c_{2}\phi_{3}(z_{1})+\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})\\ \vdots\\ \vdots\\ \phi_{N}(\eta_{N})=c_{N-2}\phi_{N-1}(z_{N-3})+\phi_{N}(z_{N-2})+\frac{c_{N-2}}{\lambda_{N}}\phi_{N}(z_{N-2})\phi_{N-1}(z_{N-3}),\\ \\ \phi_{N+1}(\eta_{N+1})=c_{N-1}\phi_{N}(z_{N-2})+\phi_{N+1}(z_{N-1})+\frac{c_{N-1}}{\lambda_{N+1}}\phi_{N+1}(z_{N-1})\phi_{N}(z_{N-2})\end{cases}$
follows at the $N+1$ step by induction. Substituting (3.165) into (3.164) yields
(3.166) $\begin{aligned} u=&u^{*}-\frac{1}{\phi_{1}(z_{N})}\left(c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}\right)\\ &+\frac{P_{1}}{\phi_{1}(z_{N})}\left(c_{N}\phi_{N}(z_{N-2})+\phi_{1}(z_{N-1})+\frac{c_{N}}{\lambda_{1}}\phi_{1}(z_{N-1})\phi_{N}(z_{N-2})\right)+\frac{P_{2}}{\phi_{1}(z_{N})}\phi_{2}(\eta_{2})\\ &+\frac{P_{3}}{\phi_{1}(z_{N})}\left(c_{1}\phi_{2}(\eta_{2})+\phi_{3}(z_{1})+\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})\right)+\frac{P_{4}}{\phi_{1}(z_{N})}\left(c_{2}\phi_{3}(z_{1})+\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})\right)\\ &+\cdots+\frac{P_{N+1}}{\phi_{1}(z_{N})}\left(c_{N-1}\phi_{N}(z_{N-2})+\phi_{N+1}(z_{N-1})+\frac{c_{N-1}}{\lambda_{N+1}}\phi_{N+1}(z_{N-1})\phi_{N}(z_{N-2})\right)\end{aligned}$
Replacing $P_{1},P_{2},\ldots,P_{N+1}$ by their respective expressions yields equation
(3.167) $\begin{aligned} u=&u^{*}-\frac{1}{\phi_{1}(z_{N})}\left(c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}\right)\\ &+\frac{\left(\phi_{1}(z_{N})-\theta\phi_{N+1}(z_{N-1})+\phi_{1}(z_{N-1})\right)}{\phi_{1}(z_{N})}\left(c_{N}\phi_{N}(z_{N-2})+\phi_{1}(z_{N-1})+\frac{c_{N}}{\lambda_{1}}\phi_{1}(z_{N-1})\phi_{N}(z_{N-2})\right)-\phi_{2}(\eta_{2})\\ &+\frac{\left(\phi_{1}(z_{N-1})-\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N})\right)}{\phi_{1}(z_{N})}\left(c_{1}\phi_{2}(\eta_{2})+\phi_{3}(z_{1})+\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})\right)\\ &+\frac{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})+\phi_{1}(z_{N})\right)}{\phi_{1}(z_{N})}\left(c_{2}\phi_{3}(z_{1})+\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})\right)\\ &+\cdots+\frac{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})+\phi_{1}(z_{N})\right)}{\phi_{1}(z_{N})}\left(c_{N-1}\phi_{N}(z_{N-2})+\phi_{N+1}(z_{N-1})+\frac{c_{N-1}}{\lambda_{N+1}}\phi_{N+1}(z_{N-1})\phi_{N}(z_{N-2})\right)\end{aligned}$
(3.168) $\begin{aligned} u=&u^{*}+c_{N}\phi_{N}(z_{N-2})+\phi_{1}(z_{N-1})+\frac{c_{N}}{\lambda_{1}}\phi_{1}(z_{N-1})\phi_{N}(z_{N-2})-\phi_{2}(\eta_{2})-c_{1}\phi_{2}(\eta_{2})-\phi_{3}(z_{1})-\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})\\ &+c_{2}\phi_{3}(z_{1})+\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})+c_{N-1}\phi_{N}(z_{N-2})+\phi_{N+1}(z_{N-1})+\frac{c_{N-1}}{\lambda_{N+1}}\phi_{N+1}(z_{N-1})\phi_{N}(z_{N-2})\\ &-\frac{1}{\phi_{1}(z_{N})}\left(c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}\right)\\ &+\frac{\left(\phi_{1}(z_{N-1})-\theta\phi_{N+1}(z_{N-1})\right)}{\phi_{1}(z_{N})}\left(c_{N}\phi_{N}(z_{N-2})+\phi_{1}(z_{N-1})+\frac{c_{N}}{\lambda_{1}}\phi_{1}(z_{N-1})\phi_{N}(z_{N-2})\right)\\ &+\frac{\left(\phi_{1}(z_{N-1})-\theta\phi_{N+1}(z_{N-1})\right)}{\phi_{1}(z_{N})}\left(c_{1}\phi_{2}(\eta_{2})+\phi_{3}(z_{1})+\frac{c_{1}}{\lambda_{3}}\phi_{3}(z_{1})\phi_{2}(\eta_{2})\right)\\ &+\frac{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)}{\phi_{1}(z_{N})}\left(c_{2}\phi_{3}(z_{1})+\phi_{4}(z_{2})+\frac{c_{2}}{\lambda_{4}}\phi_{3}(z_{1})\phi_{4}(z_{2})\right)\\ &+\cdots+\frac{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)}{\phi_{1}(z_{N})}\left(c_{N-1}\phi_{N}(z_{N-2})+\phi_{N+1}(z_{N-1})+\frac{c_{N-1}}{\lambda_{N+1}}\phi_{N+1}(z_{N-1})\phi_{N}(z_{N-2})\right).\end{aligned}$
Step 2 ($\psi_{i}\neq 0$ under Assumption H6) : Let us find the control $u$ such that the following equality is satisfied
(3.169) $\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)\dot{z}_{N-1}+\phi_{1}(z_{N})\dot{z}_{N}=c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}+\sum_{i=1}^{N+1}A_{i}\,\bigl{\lvert}\hat{\phi}_{i}-\phi_{i}(\eta_{i})\bigr{\rvert}$
We get from (3.146) for $N+1$ species
(3.170) $\begin{cases}\dot{z}_{N}=u^{*}-u-\hat{\phi}_{2}+\hat{\phi}_{1}+\hat{\phi}_{N+1}+\hat{\phi}_{N}+\cdots+\hat{\phi}_{4}-\hat{\phi}_{3},\\ \dot{z}_{N-1}=-\hat{\phi}_{1}+\hat{\phi}_{N+1}+\hat{\phi}_{N}+\cdots+\hat{\phi}_{4}-\hat{\phi}_{3}.\end{cases}$
Thus,
(3.171) $\begin{aligned} \left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)\dot{z}_{N-1}+\phi_{1}(z_{N})\dot{z}_{N}=&\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)\left(-\hat{\phi}_{1}+\hat{\phi}_{N+1}+\hat{\phi}_{N}+\cdots+\hat{\phi}_{4}-\hat{\phi}_{3}\right)\\ &+\phi_{1}(z_{N})\left(-\hat{\phi}_{2}+\hat{\phi}_{1}+\hat{\phi}_{N+1}+\hat{\phi}_{N}+\cdots+\hat{\phi}_{4}-\hat{\phi}_{3}\right)+\phi_{1}(z_{N})(u^{*}-u),\end{aligned}$
(3.172) $\begin{aligned} \left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)\dot{z}_{N-1}+\phi_{1}(z_{N})\dot{z}_{N}=&\underbrace{\left(\phi_{1}(z_{N})-\theta\phi_{N+1}(z_{N-1})+\phi_{1}(z_{N-1})\right)}_{A_{1}}\hat{\phi}_{1}\underbrace{-\phi_{1}(z_{N})}_{A_{2}}\hat{\phi}_{2}\\ &+\underbrace{\left(\phi_{1}(z_{N-1})-\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N})\right)}_{A_{3}}\hat{\phi}_{3}\\ &+\underbrace{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})+\phi_{1}(z_{N})\right)}_{A_{4}}\hat{\phi}_{4}\\ &+\cdots\cdots+\underbrace{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})+\phi_{1}(z_{N})\right)}_{A_{N+1}}\hat{\phi}_{N+1}+\phi_{1}(z_{N})(u^{*}-u)\end{aligned}$
(3.173) $\begin{aligned} \left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})\right)\dot{z}_{N-1}+\phi_{1}(z_{N})\dot{z}_{N}=&\sum_{i=1}^{N+1}A_{i}|\hat{\phi}_{i}-\phi_{i}(\eta_{i})|+\underbrace{\left(\phi_{1}(z_{N})-\theta\phi_{N+1}(z_{N-1})+\phi_{1}(z_{N-1})\right)}_{A_{1}}\phi_{1}(\eta_{1})\underbrace{-\phi_{1}(z_{N})}_{A_{2}}\phi_{2}(\eta_{2})\\ &+\underbrace{\left(\phi_{1}(z_{N-1})-\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N})\right)}_{A_{3}}\phi_{3}(\eta_{3})\\ &+\underbrace{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})+\phi_{1}(z_{N})\right)}_{A_{4}}\phi_{4}(\eta_{4})\\ &+\cdots\cdots+\underbrace{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})+\phi_{1}(z_{N})\right)}_{A_{N+1}}\phi_{N+1}(\eta_{N+1})+\phi_{1}(z_{N})(u^{*}-u).\end{aligned}$
Equation (3.157) yields the following equality
(3.174) $\begin{aligned} c_{N}\,\phi_{1}^{2}\bigl{(}z_{N-1}\bigr{)}-\theta\,c_{N}\,\phi_{N+1}^{2}\bigl{(}z_{N-1}\bigr{)}-c_{N+1}\phi_{1}^{2}\bigl{(}z_{N}\bigr{)}=&\underbrace{\left(\phi_{1}(z_{N})-\theta\phi_{N+1}(z_{N-1})+\phi_{1}(z_{N-1})\right)}_{A_{1}}\phi_{1}(\eta_{1})\\ &\underbrace{-\phi_{1}(z_{N})}_{A_{2}}\phi_{2}(\eta_{2})+\underbrace{\left(\phi_{1}(z_{N-1})-\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N})\right)}_{A_{3}}\phi_{3}(\eta_{3})\\ &+\underbrace{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})+\phi_{1}(z_{N})\right)}_{A_{4}}\phi_{4}(\eta_{4})\\ &\qquad\qquad\qquad\qquad\vdots\\ &+\underbrace{\left(\theta\phi_{N+1}(z_{N-1})-\phi_{1}(z_{N-1})+\phi_{1}(z_{N})\right)}_{A_{N+1}}\phi_{N+1}(\eta_{N+1})\\ &+\phi_{1}(z_{N})(u^{*}-u).\end{aligned}$
The goal is to find the control $u$ satisfying (3.150). By straightforward calculations, we derive the control $u$ shown in equation (3.164). ∎
Remark 3.20.
In the case of three species, we have $c_{1}=\frac{\lambda_{3}}{\lambda_{2}},\;c_{2}=\frac{\lambda_{1}}{\lambda_{3}}$. For the four-species case, $c_{1}=\frac{\lambda_{3}}{\lambda_{2}},\;c_{2}=\frac{\lambda_{4}}{\lambda_{3}},\;c_{3}=\frac{\lambda_{1}}{\lambda_{4}}$. By construction, for $N+1$ species we obtain $c_{1}=\frac{\lambda_{3}}{\lambda_{2}},\;c_{2}=\frac{\lambda_{4}}{\lambda_{3}},\;c_{3}=\frac{\lambda_{5}}{\lambda_{4}},\cdots,c_{N-1}=\frac{\lambda_{N+1}}{\lambda_{N}},\;c_{N}=\frac{\lambda_{1}}{\lambda_{N+1}}$.
We therefore obtain the following result.
Theorem 3.21.
There exists a feedback control law under which the general non-transitive competition system (3.17) is globally asymptotically stable. Furthermore, the feedback control constructed remains uniformly bounded. In particular, the control $u$ defined in (3.164) and the Lyapunov function $V_{N+1}$ in (3.154) ensures the global asymptotic stability of the $N+1$-species system (3.146). Moreover, the control satisfies $u(t)>0$ for every $t>0,$ for every $\eta_{i}(0)$ belonging to the largest level set of $V_{N+1}(\eta_{2},z_{1},\cdots,z_{N})$ within the set
[…]

### v2

@@ 3.3. Toward generalization
    (3.132) $\begin{aligned} \dot{V}_{N}(\eta_{2},z^{N}_{1},\cdots,z^{N}_{N-1})=&\phi_{1}(\eta_{1})\left(\phi_{1}(z^{N}_{N-1})-\theta\phi_{N}(z^{N}_{N-2})\right)-\phi_{1}(z^{N}_{N-1})\phi_{2}(\eta_{2})\\ &+\sum_{k=3}^{N}\left(\displaystyle\theta\sum_{i=\max(1,k-3)}^{N-2}[(-1)^{\,i+\varepsilon_{k}}\,\phi_{i+2}(z^{N}_{i})+(-1)^{\,N+1-\varepsilon_{k}}\,\phi_{1}(z^{N}_{N-1})-\delta_{k3}\theta\phi_{2}(\eta_{2})]\right)\,\phi_{k}(\eta_{k})+\phi_{1}(z^{N}_{N-1})(u-u^{*}).\end{aligned}$
    Hence, by choosing a general control $u$ of the form (3.127), we obtain relation (3.125).
Remark 3.22.
In the case of three species, we have
(3.133) $c_{1}=\frac{\lambda_{3}}{\lambda_{2}},\;c_{2}=\frac{\lambda_{1}}{\lambda_{3}},$
for the four-species case,
(3.134) $c_{1}=\frac{\lambda_{3}}{\lambda_{2}},\;c_{2}=\frac{\lambda_{4}}{\lambda_{3}},\;c_{3}=\frac{\lambda_{1}}{\lambda_{4}}.$
By construction, for $N$ species we obtain
(3.135) $c_{1}=\frac{\lambda_{3}}{\lambda_{2}},\;c_{2}=\frac{\lambda_{4}}{\lambda_{3}},\;c_{3}=\frac{\lambda_{5}}{\lambda_{4}},\cdots,c_{N-2}=\frac{\lambda_{N}}{\lambda_{N-1}},\;c_{N-1}=\frac{\lambda_{1}}{\lambda_{N}}.$
The choice of the coefficients $c_{i}$ plays a central role in preserving the structure and topology of the model while preventing the cancellation of cross terms in the context of non-transitive competition. Such cancellations remove essential information about interspecific interactions, driving the system toward a collection of decoupled subsystems, which distorts the dynamical analysis and conceals key model mechanisms (e.g., cyclic dominance). Furthermore, these cancellations suppress cross dissipation and restrict the propagation of the reduced control effect. The control $u$ defined in (3.127) provides a general control law that ensures global asymptotic stability of the system for any number of species $N\geq 3$ under non-transitive competition.
Lemma 3.23.
Let $\lambda_{1}>0$ and $\lambda_{N}>0$ and consider the closed-loop dynamics obtained via the backstepping procedure. Assume the Lyapunov weights satisfy $c_{N-1}=\lambda_{1}/\lambda_{N}>0$. Then, for any solution $(\eta(t),\psi(t))$ starting in the invariant set $\mathcal{K}_{N}$, one has
$\lambda_{1}\big(e^{z^{N}_{N-1}(t)}-1\big)\neq 0,\qquad\forall t\geq 0.$
  • Proof  :
    Step 1. Characterization of zeros. Since $\lambda_{1}>0$, the condition
    $\lambda_{1}(e^{z^{N}_{N-1}}-1)=0$
    is equivalent to $z^{N}_{N-1}=0$. Thus, it suffices to prove that $z^{N}_{N-1}(t)\neq 0$ for all $t\geq 0$.
    Step 2. Lyapunov construction and explicit constant. From the recursive backstepping design, the Lyapunov functional at stage $N$ is
    $V_{N}(\cdot)=\theta\lambda_{2}(e^{\eta_{2}}-1-\eta_{2})+\theta\sum_{i=1}^{N-2}\lambda_{i+2}(e^{z^{N}_{i}}-1-z^{N}_{i})+\lambda_{1}\big(e^{z^{N}_{N-1}}-1-z^{N}_{N-1}\big),$
    which is positive definite in $z^{N}_{N-1}$ and minimized at $z^{N}_{N-1}=0$. The cancellations in the design yield
    $\dot{V}_{N}\;=\;-\,\frac{\lambda_{1}}{\lambda_{N}}\,(e^{z^{N}_{N-1}}-1)^{2}+\text{(nonpositive terms in the other $z^{N}_{i}$)}.$
    Equivalently,
    $(e^{z^{N}_{N-1}}-1)\,\dot{z}^{N}_{N-1}\;=\;-\,\frac{\lambda_{1}}{\lambda_{N}}\,(e^{z^{N}_{N-1}}-1)^{2}.$
    On the domain where $e^{z^{N}_{N-1}}-1\neq 0$, this gives the exact closed scalar dynamics
    (3.136) $\dot{z}^{N}_{N-1}(t)\;=\;-\,\frac{\lambda_{1}}{\lambda_{N}}\,(e^{z^{N}_{N-1}(t)}-1).$
    Step 3. Sign invariance. Consider the comparison ODE
    $\dot{y}(t)=-\,\frac{\lambda_{1}}{\lambda_{N}}\,(e^{y(t)}-1).$
    If $y(0)>0$, then $\dot{y}<0$ and $y(t)$ decreases while remaining strictly positive. If $y(0)<0$, then $\dot{y}>0$ and $y(t)$ increases while remaining strictly negative. Hence the sign of $z^{N}_{N-1}(t)$ is preserved: it cannot cross zero.
    Step 4. Non-attainment of zero in finite time. Suppose, for contradiction, that $z^{N}_{N-1}(t^{\ast})=0$ for some finite $t^{\ast}$. Separating variables in (3.136), we obtain
    $\int_{z^{N}_{N-1}(0)}^{z^{N}_{N-1}(t)}\frac{1}{e^{\xi}-1}\,d\xi=-\,\frac{\lambda_{1}}{\lambda_{N}}\,t.$
    The left-hand side diverges logarithmically as $z^{N}_{N-1}(t)\rightarrow 0$, while the right-hand side remains finite for any finite $t$. This contradiction shows that $z^{N}_{N-1}(t)$ never reaches 0 in finite time.
    Starting from any $z^{N}_{N-1}(0)\neq 0$, the trajectory satisfies $z^{N}_{N-1}(t)\neq 0$ for all $t\geq 0$, and tends to 0 only asymptotically. Therefore $\lambda_{1}(e^{z^{N}_{N-1}(t)}-1)\neq 0$ for all $t\geq 0$.
Let us the following result.
Theorem 3.24.
There exists a feedback control law under which the general non-transitive competition system (3.16) is globally asymptotically stable. Furthermore, the feedback control constructed remains uniformly bounded. In particular, the control $u$ defined in (3.127) and the Lyapunov function $V_{N}$ in (3.124) ensures the global asymptotic stability of the NN-species system (3.123). Moreover, the control satisfies $u(t)>0$ for every $t>0,$ for every $\eta_{i}(0)$ belonging to the largest level set of $V_{N}(\eta_{2},z^{N}_{1},\cdots,z^{N}_{N-1})$ within the set
@@
  • Proof of Theorem 3.24 :